In a population with two alleles for a particular locus, A and a, the frequency of a is 0.3. What
would be the frequency of heterozygotes if the population is in Hardy-Weinberg equilibrium?
A. 0.48
B. 0.42
C. 0.30
D. 0.21
E. 0.09
Answer #1

Hardy Weinberg principle states that allele frequency and genotype frequency of a population remain constant in absence of any disturbing factors like mutation,immigration etc

HARDY WEINBERO PRINCIPLE The allele frequency & genotype frequency remain Oconstamy o Jenolype preequency on he generation to

PP + Pq + pq + q9 = I mawar p² + 2 pq + q ² = 1 (p+q)2 = 1 P+ 9 = 1

Given in the population

Frequency of allele a = 0.3

Frequency of A = 1- a

= 1- 0.3

= 0.7

A = 0.7

If a population is in Hardy-Weinberg equilibrium

The frequency of heterozygote will be = 2pq

= 2* 0.3*0.7

= 0.42

Hence answer is option b = 0.42

Plz upvote my answer Thank you!?